OpenStudy (anonymous):

what values for theta (0< theta < 2pi) satisfy the equation 2 cos theta + 1 = -cos theta

OpenStudy (loser66):

@Luigi0210

OpenStudy (luigi0210):

\[2\cos \theta+1=-\cos \theta\] Move all to one side\[3\cos \theta=-1\] \[\cos \theta=-\frac{ 1 }{ 3 }\]

OpenStudy (luigi0210):

I don't know if that's right..

OpenStudy (anonymous):

\[2 \cos \theta + 1 = - \cos \theta\] \[3 \cos \theta = -1\]\[\cos \theta = -\frac{ 1 }{ 3 }\]\[\theta = \cos^{-1} -\frac{ 1 }{ 3 }\] \[\cos \theta < 0 \] therefore is in the second and third quadrant Use the reference angle given by : \[\theta _{r}= \cos^{-1} \frac{ 1 }{ 3}\] Then you have two solution \[\pi - \theta _{r}\] and \[\pi + \theta _{r}\]

OpenStudy (anonymous):

so 0 and 3.14?

OpenStudy (luigi0210):

It's whatever theta is -/+ pi, like vel showed

OpenStudy (anonymous):

im confused.

OpenStudy (anonymous):

Because θr is not a special case angle like 30,45,60,90 etc you need a calculator to get the number

OpenStudy (anonymous):

Im so confused with thi questions lol sorry guys maybe because im a little sleep..

OpenStudy (anonymous):

sleepy*

OpenStudy (anonymous):

\[\cos^{-1} \frac{ 1 }{ 3 }=1.23\] \[\pi =3.14\] 3.14 - 1.23 3.14 + 1.23